# 4.2cm4.2cm3.6cm3cm

### v1

@@ 5. Automorphism groups on $\mathcal{L}(\Gamma)$
Let $L=\mathrm{span}_{\mathbb{F}}\{L_{\alpha,i}\,|\,\alpha\in\Gamma,i\in\mathbb{Z}\},\,H=\mathrm{span}_{\mathbb{F}}\{H_{\alpha,i}\,|\,\alpha\in\Gamma,i\in\mathbb{Z}\}.$
@@
$I=\left\{H_{\alpha,i}\,|\,\alpha\in\Gamma,i\in\mathbb{Z}\right\}$ is a unique maximal ideal on $\mathcal{L}(\Gamma)$.
@@
By the definition of $\mathcal{L}(\Gamma)$, $I$ is an ideal of $\mathcal{L}(\Gamma)$. Since $\mathcal{L}(\Gamma)/I$ is isomorphic to the generalized centerless Loop Virasoro algebra. The centerless Loop Virasoro algebra is a simple Lie algebra, then $I$ is a maximal ideal. Next, we prove $I=I^{\prime}$. Let $I^{\prime}$ be another maximal ideal of $\mathcal{L}(\Gamma)$. For all $0\neq x\in I^{\prime}$, let $x=l+h$, where $l\in L,\,h\in H$. If $l\neq 0$, then $I^{\prime}=\mathcal{L}(\Gamma)$, which is a contradiction. Thus $l=0$. ∎
@@
Let $\mathcal{LV}(\Gamma)$ and $\mathcal{LV}(\Gamma^{\prime})$ be two generalized centerless Loop Virasoro algebra. $\mathcal{LV}(\Gamma)\cong\mathcal{LV}(\Gamma^{\prime})$ if and only if there there exists $a\in\mathbb{F}^{\ast}$, such that $a\Gamma^{\prime}=\Gamma$, where $\mathbb{F}^{\ast}=\mathbb{F}\setminus\{0\}$.
@@
Notice that the ”only if” part is obvious. We now prove the ”if” part.
@@
$\theta[L_{0,0},L_{\alpha,j}]$ $=$ $\theta(-\alpha L_{\alpha,j})$
@@
$\theta[L_{0,0},L_{\alpha,j}]$ $=$ $[\theta(L_{0,0}),\theta(L_{\alpha,j})]$
@@
Thus $[L^{\prime}_{0,0},\theta(L_{\alpha,j})]=-a^{-1}\alpha\theta(L_{\alpha,j})$, that is $-a^{-1}\alpha\in\Gamma^{\prime}$, we conclude $\Gamma\subset a\Gamma^{\prime}$. Similarly, it can be proved that $a\Gamma^{\prime}\subset\Gamma$. ∎
@@
$\mathcal{L}(\Gamma)\cong\mathcal{L}(\Gamma^{\prime})$ if and only if there exists $a\in\mathbb{F}^{\ast}$, such that $a\Gamma^{\prime}=\Gamma$.
@@
Let $\theta:\mathcal{L}(\Gamma)\rightarrow\mathcal{L}(\Gamma^{\prime})$ be a Lie algebra isomorphism and $I,I^{\prime}$ be the maximal ideals of $\mathcal{L}(\Gamma)$ and $\mathcal{L}(\Gamma^{\prime})$, respectively. By Lemma 19, we have $\theta(I)=I^{\prime}$, and by Lemma 20, we have $\mathcal{L}(\Gamma)/I\cong\mathcal{L}(\Gamma^{\prime})/I^{\prime}$, then exists $a\in\mathbb{F}^{\ast}$.
@@
[…]
@@
For all $\theta\in\mathrm{Aut}\mathcal{L}(\Gamma)$, considering $\theta$ act on the basis of $\mathcal{L}(\Gamma)$, for all $\alpha\in\Gamma,\,i\in\mathbb{Z}$, assume $f_{\alpha,i}\in\mathbb{F}[t^{1},t^{-1}]\in H,$
@@
By comparing the expressions, we have:
@@
$a\beta f_{\alpha+\beta,i+j}=\beta f_{\alpha,i}f_{\beta,j},\,\forall\alpha,\beta\in\Gamma,\,i,j\in\mathbb{Z}.$
@@
Let $a\in\mathbb{F}^{\ast},\,f_{\alpha},f_{\beta}\in\mathbb{F}[t^{1},t^{-1}]$. For all $\alpha,\beta\in\Gamma,\,i,j\in\mathbb{Z}$, then $af_{\alpha+\beta,i+j}=f_{\alpha,i}f_{\beta,j}$.
@@
The proof is similar to that in Lemma 3 and Lemma 4. ∎
Let $\mathbb{F}[t^{1},t^{-1}]^{\ast}=\{at^{i}\,|\,a\in\mathbb{F}^{\ast},i\in\mathbb{Z}\}$. By Lemma 22, we have $f_{0,0}$ is a non-zero element on $\mathbb{F}[t^{1},t^{-1}]$, then $f_{0,0}\in\mathbb{F}[t^{1},t^{-1}]^{\ast}$. Thus, we can write $f_{\alpha,i}$ in the following form:
$f_{\alpha,i}:=a\mu(\alpha,i)t^{\epsilon(\alpha,i)},\,\forall\alpha\in\Gamma,\,i\in\mathbb{Z},$
where $\mu(\alpha,i)\in\mathbb{F}^{\ast},\,\epsilon(\alpha,i)\in\mathbb{Z}$, thus for all $\alpha_{1},\alpha_{2}\in\Gamma,\,i,j\in\mathbb{Z}$,
@@
(i) For all $a\in A=\{a\in\mathbb{F}^{\ast}\,|\,a\Gamma=\Gamma\}$, the map
@@
[…]
@@
(ii) For all $\phi\in\mathrm{Hom}(\Gamma,\mathbb{Z})$, the map
@@
[…]
@@
(iii) For all $\chi\in\chi(\Gamma)$, $\chi(\Gamma)$ is the set consisting of group homomorphisms from $\Gamma$ to $\mathbb{F}^{\ast}$, the map
@@
[…]
@@
[…]
@@
[…]
@@
Recall the [7], we have the set $(A\times\mathrm{Hom}(\Gamma,\mathbb{Z})\times\chi(\Gamma)\times\mathrm{Aut}\mathbb{Z}\times\mathbb{F^{\ast}})$ is a multiplicative group for which $(1,0,0,1,1)$ is the identify element, it satisfies
@@
where $\iota_{a}$ is an automorphism on $\Gamma$, satisfying $\iota_{a}(\alpha)=a\alpha$. If $b^{\phi}\in\chi(\Gamma)$, it satisfies $(b^{\phi})(\alpha)=b^{\phi(\alpha)}$, where $\alpha\in\Gamma,\,b\in\mathbb{F}^{\ast},\,\phi\in\mathrm{Hom}(\Gamma,\mathbb{Z}).$
@@
[…]
@@
$=$ $a_{1}a_{2}(b_{1}^{\psi_{2}}b_{2})^{i}b_{1}^{\phi_{2}(\alpha)}\chi_{2}(\alpha)\chi(\frac{\alpha}{a_{2}})L_{\frac{\alpha}{a_{1}a_{2}},(\psi_{1}\psi_{2})(i)+(\phi_{1}\iota_{a^{-1}_{2}}+\psi_{1}\phi_{2})(\alpha)})$
@@
Let $\theta_{a_{1}}\theta_{\phi_{1}}\theta_{\chi_{1}}\theta_{\psi_{1}}\theta_{b_{1}}=\theta_{a_{2}}\theta_{\phi_{2}}\theta_{\chi_{2}}\theta_{\psi_{2}}\theta_{b_{2}}$. By $\theta_{a},\theta_{\phi},\theta_{\chi},\theta_{\psi},\theta_{b}$ are automorphisms on $\mathcal{L}(\Gamma)$, thus $\theta_{a_{1}}=\theta_{a_{2}},\theta_{\phi_{1}}=\theta_{\phi_{2}},\,\theta_{\chi_{1}}=\theta_{\chi_{2}},\,\theta_{\psi_{1}}=\theta_{\psi_{2}},\theta_{b_{1}}=\theta_{b_{2}}$, we conclude $(a_{1},\phi_{1},\chi_{1},\psi_{1},b_{1})=(a_{2},\phi_{2},\chi_{2},\psi_{2},b_{2})$, $\Psi$ is an injective map.
For all $\theta\in\mathrm{Aut}\mathcal{L}(\Gamma)$, by Lemma 21 and Lemma 22 we have $\theta(L_{\alpha,i})=a\mu(\alpha,i)L_{{\frac{\alpha}{a}},\epsilon(\alpha,i)},\theta(H_{\alpha,i})=a\mu(\alpha,i)H_{{\frac{\alpha}{a}},\epsilon(\alpha,i)}$. Let $\mu(\alpha,0)=\chi(\alpha),\,\mu(0,i)=b(i),\,\epsilon(\alpha,0)=\phi(\alpha),\,\epsilon(0,i)=\psi(i)$. Then $\Psi(a,\phi,\chi,\psi,b)=\theta$, therefore $\Psi$ is a surjective map. In conclusion, $\Psi$ is an isomorphism. ∎

### v2

@@ 5. Automorphism groups on $\mathcal{L}(\Gamma)$
Let $L=\mathrm{span}\{L_{\alpha,i}\,|\,\alpha\in\Gamma,i\in\mathbb{Z}\},\,H=\mathrm{span}\{H_{\alpha,i}\,|\,\alpha\in\Gamma,i\in\mathbb{Z}\}.$
@@
$H$ is a unique maximal ideal on $\mathcal{L}(\Gamma)$.
@@
By the definition of $\mathcal{L}(\Gamma)$, $H$ is an ideal of $\mathcal{L}(\Gamma)$. Since $\mathcal{L}(\Gamma)/H$ is isomorphic to the generalized centerless loop Virasoro algebra. The centerless loop Virasoro algebra is a simple Lie algebra, then $H$ is a maximal ideal. Next, we prove $H=I$. Let $I$ be another maximal ideal of $\mathcal{L}(\Gamma)$. For all $0\neq x\in I$, let $x=l+h$, where $l\in L,\,h\in H$. If $l\neq 0$, then $I=\mathcal{L}(\Gamma)$, which is a contradiction. Thus $l=0$. ∎
@@
Let $\mathcal{LV}(\Gamma)$ and $\mathcal{LV}(\Gamma^{\prime})$ be two generalized centerless loop Virasoro algebra. Then $\mathcal{LV}(\Gamma)\cong\mathcal{LV}(\Gamma^{\prime})$ if and only if there exists $a\in\mathbb{F}^{\ast}$ such that $a\Gamma^{\prime}=\Gamma$, where $\mathbb{F}^{\ast}=\mathbb{F}\setminus\{0\}$.
@@
Notice that the “only if” part is obvious. We now prove the “if” part.
@@
$\theta([L_{0,0},L_{\alpha,j}])$ $=$ $\theta(-\alpha L_{\alpha,j})$
@@
$\theta([L_{0,0},L_{\alpha,j}])$ $=$ $[\theta(L_{0,0}),\theta(L_{\alpha,j})]$
@@
Thus $L^{\prime}_{0,0},\theta(L_{\alpha,j})]=-a^{-1}\alpha\theta(L_{\alpha,j})$, that is $a^{-1}\alpha\in\Gamma^{\prime}$, we conclude $\Gamma\subset a\Gamma^{\prime}$. Similarly, it can be proved that $a\Gamma^{\prime}\subset\Gamma$. ∎
@@
$\mathcal{L}(\Gamma)\cong\mathcal{L}(\Gamma^{\prime})$ if and only if there exists $a\in\mathbb{F}^{\ast}$ such that $a\Gamma^{\prime}=\Gamma$.
@@
Let $\theta:\mathcal{L}(\Gamma)\rightarrow\mathcal{L}(\Gamma^{\prime})$ be a Lie algebra isomorphism and $I,I^{\prime}$ be the maximal ideals of $\mathcal{L}(\Gamma)$ and $\mathcal{L}(\Gamma^{\prime})$, respectively. By Lemma 19, we have $\theta(I)=I^{\prime}$, and by Lemma 20, we have $\mathcal{L}(\Gamma)/I\cong\mathcal{L}(\Gamma^{\prime})/I^{\prime}$, then there exists $a\in\mathbb{F}^{\ast}$ such that $a\Gamma^{\prime}=\Gamma$.
@@
$L_{\alpha,i}$ $\longmapsto$ $L_{\frac{\alpha}{a},i}$
$H_{\alpha,i}$ $\longmapsto$ $H_{\frac{\alpha}{a},i},\,\alpha\in\Gamma,\,i\in\mathbb{Z}.$
@@
For all $\theta\in\mathrm{Aut}\mathcal{L}(\Gamma)$, considering $\theta$ act on the basis of $\mathcal{L}(\Gamma)$ for all $\alpha\in\Gamma,\,i\in\mathbb{Z}$, assume $f_{\alpha,i}\in\mathbb{F}[t,t^{-1}],$
@@
By comparing the above expressions, we have:
@@
$a\beta f_{\alpha+\beta,i+j}=\beta f_{\alpha,i}f_{\beta,j},\,\mbox{for all}\,\alpha,\beta\in\Gamma\,\mbox{and all}\,i,j\in\mathbb{Z}.$
@@
Let $a\in\mathbb{F}^{\ast},\,f_{\alpha,i},f_{\beta,j}\in\mathbb{F}[t,t^{-1}]$. Then $af_{\alpha+\beta,i+j}=f_{\alpha,i}f_{\beta,j}$ for all $\alpha,\beta\in\Gamma$ and all $i,j\in\mathbb{Z}$.
@@
The proof is similar with that in Lemma 3 and Lemma 4. ∎
Let $\mathbb{F}[t,t^{-1}]^{\ast}=\{at^{i}\,|\,a\in\mathbb{F}^{\ast},i\in\mathbb{Z}\}$. By Lemma 22, we have $f_{0,0}$ is a non-zero element on $\mathbb{F}[t,t^{-1}]$, then $f_{0,0}\in\mathbb{F}[t,t^{-1}]^{\ast}$. Thus, we can write $f_{\alpha,i}$ in the following form:
$f_{\alpha,i}:=a\mu(\alpha,i)t^{\epsilon(\alpha,i)},\,\mbox{for all}\,\alpha\in\Gamma\,\mbox{and all}\,i\in\mathbb{Z},$
where $\mu(\alpha,i)\in\mathbb{F}^{\ast},\,\epsilon(\alpha,i)\in\mathbb{Z}$, thus for all $\alpha_{1},\alpha_{2}\in\Gamma$ and all $i,j\in\mathbb{Z}$,
@@
(i) For all $a\in A=\{a\in\mathbb{F}^{\ast}\,|\,a\Gamma=\Gamma\}$, the map is given by
@@
$L_{\alpha,i}$ $\longmapsto$ $aL_{\frac{\alpha}{a},i}$
$H_{\alpha,i}$ $\longmapsto$ $aH_{\frac{\alpha}{a},i},\,\mbox{for all}\,\alpha\in\Gamma,\,\mbox{and all}\,i\in\mathbb{Z}$
@@
(ii) For all $\phi\in\mathrm{Hom}(\Gamma,\mathbb{Z})$, the map is given by
@@
$L_{\alpha,i}$ $\longmapsto$ $L_{\alpha,i+\phi(\alpha)}$
$H_{\alpha,i}$ $\longmapsto$ $H_{\alpha,i+\phi(\alpha)},\,\mbox{for all}\,\alpha\in\Gamma,\,\mbox{and all}\,i\in\mathbb{Z}$
@@
(iii) For all $\chi\in\chi(\Gamma)$, $\chi(\Gamma)$ is the set consisting of group homomorphisms from $\Gamma$ to $\mathbb{F}^{\ast}$, the map is given by
@@
$L_{\alpha,i}$ $\longmapsto$ $\chi(\alpha)L_{\alpha,i}$
$H_{\alpha,i}$ $\longmapsto$ $\chi(\alpha)H_{\alpha,i},\,\mbox{for all}\,\alpha\in\Gamma,\,\mbox{and all}\,i\in\mathbb{Z}$
@@
$L_{\alpha,i}$ $\longmapsto$ $L_{\alpha,\psi(i)}$
$H_{\alpha,i}$ $\longmapsto$ $H_{\alpha,\psi(i)},\,\mbox{for all}\,\alpha\in\Gamma,\,\mbox{and all}\,i\in\mathbb{Z}$
@@
$L_{\alpha,i}$ $\longmapsto$ $b^{i}L_{\alpha,i}$
$H_{\alpha,i}$ $\longmapsto$ $b^{i}H_{\alpha,i},\,\mbox{for all}\,\alpha\in\Gamma,\,\mbox{and all}\,i\in\mathbb{Z}$
@@
Recall the [19], we have the set $(A\times\mathrm{Hom}(\Gamma,\mathbb{Z})\times\chi(\Gamma)\times\mathrm{Aut}\mathbb{Z}\times\mathbb{F^{\ast}})$ is a multiplicative group for which $(1,0,0,1,1)$ is the identity element, it satisfies
@@
where $\iota_{a}$ is an automorphism on $\Gamma$, satisfying $\iota_{a}(\alpha)=a\alpha$. If $b^{\phi}\in\chi(\Gamma)$, then $(b^{\phi})(\alpha)=b^{\phi(\alpha)}$, where $\alpha\in\Gamma,\,b\in\mathbb{F}^{\ast},\,\phi\in\mathrm{Hom}(\Gamma,\mathbb{Z}).$
@@
$(a,\phi,\chi,\psi,b)$ $\longmapsto$ $\theta_{a}\theta_{\phi}\theta_{\chi}\theta_{\psi}\theta_{b}$
@@
$=$ $a_{1}a_{2}(b_{1}^{\psi_{2}}b_{2})^{i}b_{1}^{\phi_{2}(\alpha)}\chi_{2}(\alpha)\chi(\frac{\alpha}{a_{2}})L_{\frac{\alpha}{a_{1}a_{2}},(\psi_{1}\psi_{2})(i)+(\phi_{1}\iota_{a^{-1}_{2}}+\psi_{1}\phi_{2})(\alpha)}$
@@
Let $\theta_{a_{1}}\theta_{\phi_{1}}\theta_{\chi_{1}}\theta_{\psi_{1}}\theta_{b_{1}}=\theta_{a_{2}}\theta_{\phi_{2}}\theta_{\chi_{2}}\theta_{\psi_{2}}\theta_{b_{2}}$. Since $\theta_{a},\theta_{\phi},\theta_{\chi},\theta_{\psi},\theta_{b}$ are automorphisms on $\mathcal{L}(\Gamma)$, then $\theta_{a_{1}}=\theta_{a_{2}},\theta_{\phi_{1}}=\theta_{\phi_{2}},\,\theta_{\chi_{1}}=\theta_{\chi_{2}},\,\theta_{\psi_{1}}=\theta_{\psi_{2}},\theta_{b_{1}}=\theta_{b_{2}}$,furthermore, we conclude $(a_{1},\phi_{1},\chi_{1},\psi_{1},b_{1})=(a_{2},\phi_{2},\chi_{2},\psi_{2},b_{2})$, $\Psi$ is an injective map.
For all $\theta\in\mathrm{Aut}\mathcal{L}(\Gamma)$, by Lemma 21 and Lemma 22, we have $\theta(L_{\alpha,i})=a\mu(\alpha,i)L_{{\frac{\alpha}{a}},\epsilon(\alpha,i)},\theta(H_{\alpha,i})=a\mu(\alpha,i)H_{{\frac{\alpha}{a}},\epsilon(\alpha,i)}$. Let $\mu(\alpha,0)=\chi(\alpha),\,\mu(0,i)=b(i),\,\epsilon(\alpha,0)=\phi(\alpha),\,\epsilon(0,i)=\psi(i)$. Then $\Psi(a,\phi,\chi,\psi,b)=\theta$, therefore $\Psi$ is a surjective map. In conclusion, $\Psi$ is an isomorphism. ∎